OpenStudy (anonymous):

WRITE AN EQUATION IN SLOPE-INTERCEPT, POINT-SLOPE, OR STANDARD FORM FOR THE LINE WITH THE GIVEN INFORMATION. EXPLAIN WHY YOU CHOSE THE FORM YOU USED. a. passes through (-1, 4) and (-5, 2) (standard form) b. slope 2, y-intercept -4 (slope-intercept) c. has an x-intercept of 6 and a y-intercept of 3 (standard form) d. passes through (1, 2) with slope -5/3 (point-slope)

OpenStudy (anonymous):

I figured out which form they're in by common knowledge and I know how to explain WHY I choose each form, but I don't know how to write and equation in each form, and I mainly need help with that. Please and thank you :D Medals will be given to best/most helpful answer(s).

OpenStudy (anonymous):

Point-Slope Formula \((y-y_1)=m(x-x_1)\) Slope-Intercept Formula \(y=mx+b\)

OpenStudy (anonymous):

so for the point-slope equation I would put (2 - -5) = m(1 - 3)? Is that correct? And for slope-intercept equation I would put -4 = m2 + b? Is that correct? I'm sorry if I'm being difficult I'm just not so good with these sort of things.

OpenStudy (anonymous):

On the Point-Slope, you will choose a point and substitute the values for x and y in place of \(x_1\) and \(y_1\). Also, find the slope and plug that in for m.

OpenStudy (anonymous):

Do I choose any slope? And how would I find the slope?

OpenStudy (anonymous):

I meant any point. Oops :3

OpenStudy (anonymous):

Slope is the change in \(y\) over the change in \(x\). The formula looks like this: \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

And yes, you can choose any point on the line to create the equation.

OpenStudy (anonymous):

Okay I understand how to use the formula, but what would I input into it to find the slope?

OpenStudy (anonymous):

Line? Does this mean I have to draw a graph in order to find my point?

OpenStudy (anonymous):

Point 1 (-1,4) Point 2 (-5,2) Slope: \(m=\frac{2-4}{-5-(-1)}\) No, you do not have to draw a graph, just choose one of the given points.

OpenStudy (anonymous):

Oh okay, that makes sense lol. Thank you, I would've never figured this stuff out on my own.

OpenStudy (anonymous):

You are welcome. FYI: www.purplemath.com is also a good resource.

OpenStudy (anonymous):

Thanks a lot, so once I figured out the slope with the formula, take one of the given points and insert them in the equation?

OpenStudy (anonymous):

Exactly! :-)

OpenStudy (anonymous):

So say if I chose point 1 (-1, 4) it would be, (4 - 4) = 3.8(-1 - [-1]) ?

OpenStudy (anonymous):

Too many numbers. You still need the x and y in the equation. Only substitute one point for the \(x_1\) and \(y_1\).

OpenStudy (anonymous):

So it would be (y - 4) = 3.8(x - [-1])?

OpenStudy (anonymous):

Mostly. Where did the 3.8 come from?

OpenStudy (anonymous):

Oops I meant 0.5, I think I accidentally copied and pasted that into the equation from a different problem.

OpenStudy (anonymous):

In that case, you are correct. And that should change your choice of equation format.

OpenStudy (anonymous):

How would that change the equation format?

OpenStudy (anonymous):

You answered standards form. It has been a few years so I could be wrong, but I believe the equation you just wrote is in point-slope form.

OpenStudy (anonymous):

I thought the format you gave me was point-slope form?

OpenStudy (anonymous):

Yes, the formula I gave you is Point-Slope, but you answered Standard Form up at the top. I believe your answer should be Point-Slope.

OpenStudy (anonymous):

Oh okay I get what you were saying now, it actually gave me the correct formula. Thank you.

OpenStudy (anonymous):

Glad to help. How are you doing on the others?

OpenStudy (anonymous):

Not so good, haha.

OpenStudy (anonymous):

B should be easier - use the Slope-Intercept form since that is the information that you have.

OpenStudy (anonymous):

How do I find x-intercept? and do I just leave the variable b as is?

OpenStudy (anonymous):

Not sure why you would want the x-intercept... \{y=mx+b\) where m=slope and b= y-intercept.

OpenStudy (anonymous):

So I don't need the find x? and do I find the slope as the same way as before?

OpenStudy (anonymous):

Slope = 2, y-intercept = -4 That information was given.

OpenStudy (anonymous):

Nevermind I forgot the slope was already there

OpenStudy (anonymous):

\(:-)\)

OpenStudy (anonymous):

So it would be y = 2x + -4?

OpenStudy (anonymous):

That is the answer I got.

OpenStudy (anonymous):

Oh that's simple lol, could you help me with next one?

OpenStudy (anonymous):

c? I can help you get to the equation taking the scenic route. You would be better off reposting that question.

OpenStudy (anonymous):

Yes c and okay.

OpenStudy (anonymous):

d is similar to a with less steps.

OpenStudy (anonymous):

c. has an x-intercept of 6 and a y-intercept of 3 (I believe it's Standard Form)

OpenStudy (anonymous):

I don't know an easy way to get to the equation from there. I'm sure there is a formula or quick method, but I don't know it. Using the Point-Slope Formula will work, but I think there is an easier way.

OpenStudy (anonymous):

I'm going to sign off soon - do you want to work on d?

OpenStudy (anonymous):

Yea let's work on d

OpenStudy (anonymous):

I think I figured d out, is it (y - 2) = -1.67(x - 1)